Identify the surfaces of the following equations by converting them into equations in the Cartesian form. Show your complete solutions. z^2=4+4r^2

1
Expert's answer
2022-06-07T04:52:41-0400
"z^2=4+4r^2=>z^2=4+4x^2+4y^2""-\dfrac{x^2}{1^2}-\dfrac{y^2}{1^2}+\dfrac{z^2}{2^2}=1"

The hyperboloid of two sheets.